Determine the simplest formula of compounds whose composition are given
a)0.795g of an oxide of copper having been reduced to 0.635g of copper by heating in a stream of hydrogen
b)a compound that contains 27.38%sodium 1.19%hydogen 14.20%carbon and 57.15% oxygen
c)a compound that contains 29.41% calcium 23.53% sulphur and 47.06%oxygen
d)1.39 ofcrystalline salt that contains 0.28g iron 0.16g sulphur 0.32g oxygen and 0.63g water of crystalline
e)1.00g of a compound that contains 0.40g calcium
1
Expert's answer
2020-04-09T13:11:24-0400

a) CuxOy + H2 = Cu + H2O

0.635/0.795=0.7987=79.87

79.87/63/546=1.256

100%-79.87=20.13

79.87/63.546=20.13/x

x=16.02

CuO

b)NaxHyCuOz


"\frac{27.38}{22.98977}=\frac{1.19}{1.00794}=\frac{14.20}{12.011}=\frac{57.15}{15.9994}=1.1909645=1.18062583=1.1822496=3.5720895=1:1:1:3"

NaHCO3 sodium bicarbonate

c)СaxSyOz


"\frac{29.41}{40.08}=\frac{23.53}{32.06}=\frac{47.26}{15.9994}=0.7337=0.7339=2.95386077=1:1:4"


СaSO4 calcium sulfate

d) 0.28/1.39=20.14

0.16/1.39=11.51

0.32/1.39=23.02

0.63/1.39=45.33

"\frac{20.14}{55.847}=\frac{11.51}{32.06}=\frac{23.02}{15.994}=\frac{45.33}{18.01528}=0.3606=0.3590=1.439=2.5162=1:1:4:7"


FeSO4*7H2O Iron sulfate (II)


e)0.40/1.0=0.4=40% 100-40=60

40/40.08=60/x

x=60.12

60.12+40.08=100.2

CaCO3 - calcium carbonate